A sample of nitrogen gas expands in volume from 1.6 L to 5.4 L at constant temperature. Calculate the work done in joules if the gas expands (a) against a vacuum, (b) against a constant pressure of 0.80 atm, and (c) against a constant pressure of 3.7 atm.


1
Expert's answer
2021-09-18T23:58:12-0400

Solution


a.) "W = -0 \u00d7(5.4-1.6) L = 0" J


b.) "P = 0.80"


"V2-V1 = ( 5.4-1.6) = 3.8 L"


"1 L.atm = 101.325 J"


"W= - 0.80\u00d73.8 = -3.04"


"= -3.04\u00d7101.325 J\n\n\n\n = -308.03 J"


c.)"P = 3.7 atm"


"V2-V1 = ( 5.4-1.6) = 3.8 L"


"W = - 3.7\u00d73.8 = -14.06 L.atm"


="-14.06\u00d7101.325 J"


"= -1424.63 J"